# Exceptional supercuspidal representations in small residue characteristic

### v1

@@ 4.1 $G=G_{2}$, $q=2$
As $G(k)_{x,+}=G(k)_{x,0}$, it suffices to check that the only element $w$ of the affine Weyl group that intertwines $\chi_{\lambda_{i}}$ is trivial. By Lemma 3.18,
@@
and so $w$ is either trivial or the simple reflection $w_{\alpha_{1}}$ in $\alpha_{1}=0$. The latter sends $\alpha_{2}$ to $3\alpha_{1}+\alpha_{2}\in\Psi_{x}^{+}$, violating the condition from Lemma 3.18. Therefore $w$ must be trivial. Lemma 3.2 implies that the $\pi_{i}$ are irreducible and supercuspidal for $i=1,2$. ∎
@@ 4.2 $G=B_{n}$, $q=2$
These inequalities imply $r_{i}=0$ for all $i$, as required. Then, by Theorem 3.17, letting $\chi$ be the unique nontrivial additive character $\chi:\mathbb{F}_{2}\to\mathbb{C}^{\times}$, the compactly induced representation $\pi_{\lambda}:=\mathrm{ind}_{G(k)_{x,+}}^{G(k)}\chi_{\lambda}$ is a finite direct sum of irreducible supercuspidal representations of $G(k)$.
Remark 4.3.
@@
By the proof of Proposition 3.10, if $\pi_{\lambda}$ has a nontrivial fixed vector under $G(k)_{y^{\prime},r+}$ for some $y^{\prime}\in\mathcal{B}$, then there exists $y\in\mathcal{A}$ and $g\in\mathbb{G}_{x}(\mathbb{F}_{q})$ such that $g\cdot\lambda$ is trivial on $G(k)_{x,+}\cap G(k)_{y,r+}$. In our case $\mathbb{G}_{x}(\mathbb{F}_{2})$ is trivial. Let $y=\sum_{i}r_{i}\omega_{i}^{\vee}$ for $r_{i}\in\mathbb{R}$. Then $\lambda$ is trivial on $G(k)_{x,+}\cap G(k)_{y,r+}$ if and only if the following inequalities hold:
@@
Combining [RY14, Theorem 4.1] and [RLYG12, Section 8.2.2], the points $x\in\mathcal{A}$ at which $V_{x}^{\vee}$ admits an $\mathbb{F}_{q}$-rational stable functional satisfy $\delta(x)=m/2n$, where $m\mid n$. By Proposition 3.10, the corresponding supercuspidal representations constructed by [RY14, Proposition 2.4] have depth $m/2n$. When $n\geq 3$ this does not take the value $\frac{1}{2(n-1)}$. ∎
@@
As we are working in the setting $q=2$, we have that $G(k)_{x,+}=G(k)_{x,0}$. Note that in the simply-connected case, the $G(k)_{x,+}$-double cosets of $1$ and $\sigma$ agree, and also of $w_{\alpha_{n}}$ and $\sigma w_{\alpha_{n}}$ agree. By the affine Bruhat decomposition 3.4, it suffices to show that the elements of $W(G)$ that intertwine $\chi_{\lambda}$ are $1,w_{\alpha_{n}},\sigma$ and $\sigma w_{\alpha_{n}}$. These certainly intertwine as their action on $\Psi$ takes $\mathrm{supp}(\lambda)$ to itself so that $\lambda=\lambda^{w}$ on $G(k)_{x,+}\cap G(k)_{x,+}^{w}$ for any of $w=1,w_{\alpha_{n}},\sigma$ and $\sigma w_{\alpha_{n}}$.
On the other hand, by Lemma 3.18, if $w\in W(G)$ intertwines $\chi_{\lambda}$, then
@@
Moreover, by Lemma 3.19, the set $w(\mathcal{S}(x))\cap\mathcal{S}(x)$ has no zeros in $\overline{\mathcal{C}}$. This set must therefore contain $1-\alpha_{0},\alpha_{1},\dots,\alpha_{n-2}$ and also $\{\alpha_{n-1}+\alpha_{n}\}$, $\{\alpha_{n-1},\alpha_{n}\}$, $\{\alpha_{n-1},\alpha_{n-1}+2\alpha_{n}\}$ or $\{\alpha_{n},\alpha_{n-1}+2\alpha_{n}\}$. The only two elements of $\mathcal{S}(x)$ whose gradients pair negatively with the gradients of at least three other elements of $\mathcal{S}(x)$ are $\alpha_{2}$ and $\alpha_{n-2}$. For $\alpha_{2}$, these three elements are $1-\alpha_{0},\alpha_{1}$ and $\alpha_{3}$. As these are all in $w(\mathcal{S}(x))\cap\mathcal{S}(x)$ (as $n\geq 5$), it follows that $w(\alpha_{2})=\alpha_{2}$ or $w(\alpha_{n-2})=\alpha_{2}$. Here we use the fact that the Weyl group acts by isometries on the roots.
@@
Given $w(\alpha_{2})=\alpha_{2}$, again using combinatorics of the affine Dynkin diagram, $w$ must fix $\alpha_{3},\dots,\alpha_{n-2}$ and either fix or swap $1-\alpha_{0},\alpha_{1}$. Composing with $\sigma$, which swaps $1-\alpha_{0},\alpha_{1}$, we may assume that $w$ fixes them. We must show that $w=1$ or $w=w_{\alpha_{n}}$. Out of the remaining elements $\alpha_{n-1}$, $\alpha_{n}$, $\alpha_{n-1}+\alpha_{n}$ and $\alpha_{n-1}=2\alpha_{n}$, the only one to pair with $\alpha_{n-2}$ to 0 is $\alpha_{n}$. So if $\alpha_{n}\in w(\mathcal{S}(x))\cap\mathcal{S}(x)$ then $w(\alpha_{n})=\alpha_{n}$ and the only possibility is $w=1$. Similarly, $\alpha_{n-1}+\alpha_{n}$ is the only short root of the four that pairs negatively with $\alpha_{n-2}$. If $\alpha_{n-1}+\alpha_{n}\in w(\mathcal{S}(x))\cap\mathcal{S}(x)$ then $w(\alpha_{n-1}+\alpha_{n})=\alpha_{n-1}+\alpha_{n}$. One of $\alpha_{n-1}$ or $\alpha_{n-1}+2\alpha_{n}$ must then map under $w$ to a positive affine root at $x$, necessarily in $\mathcal{S}(x)$ by Lemma 3.18. The only possibility is that
@@ 4.3 $G=G_{2}$, $p=3$
with $\lambda_{-\dot{\psi}}\in V_{\psi}^{\vee}$. Set $\lambda_{\alpha_{0}-\alpha_{2}}=0,\lambda_{-\alpha_{1}}=0$. Take $\lambda_{\alpha_{0}}$, $\lambda_{-(\alpha_{1}+\alpha_{2})}$, $\lambda_{-(3\alpha_{1}+\alpha_{2})}$ and $\lambda_{-(3\alpha_{1}+2\alpha_{2})}$ to be any nonzero $\mathbb{F}_{q}$-linear functionals in the respective $V_{\psi}^{\vee}$. We claim that this gives an $\mathbb{F}_{q}$-stable functional.
Let $g\in\mathbb{G}_{x}(\mathbb{F}_{q})$. The factor $\mathbb{G}_{m}$ of $\mathrm{SL}_{2}\times\mathbb{G}_{m}$ acts by nonzero scaling on each summand of $V(x)^{\vee}$. Identify $V_{1-\alpha_{0}}^{\vee}\oplus V_{1-\alpha_{0}+\alpha_{2}}^{\vee}$, $V_{\alpha_{1}}^{\vee}\oplus V_{\alpha_{1}+\alpha_{2}}^{\vee}$ and $V_{3\alpha_{1}+\alpha_{2}}^{\vee}\oplus V_{3\alpha_{1}+2\alpha_{2}}^{\vee}$ as the standard representation of $\mathrm{SL}_{2}$. As $(\lambda_{\alpha_{0}},0)$, $(0,\lambda_{-(\alpha_{1}+\alpha_{2})})$ and $(\lambda_{-(3\alpha_{1}+\alpha_{2})},\lambda_{-(3\alpha_{1}+2\alpha_{2})})$ are pairwise linearly independent under this identification, so are their images under the action of $g$. Therefore, $\mathrm{supp}(g\cdot\lambda)$ contains at least two elements of each of $\{1-\alpha_{0},\alpha_{1},3\alpha_{1}+\alpha_{2}\}$ and $\{1-\alpha_{0}+\alpha_{2},\alpha_{1}+\alpha_{2},3\alpha_{1}+2\alpha_{2}\}$. It also contains at least one element of each of $\{1-\alpha_{0},1-\alpha_{0}+\alpha_{2}\}$, $\{\alpha_{1},\alpha_{1}+\alpha_{2}\}$ and $\{3\alpha_{1}+\alpha_{2},3\alpha_{1}+2\alpha_{2}\}$.
@@
$\{1-\alpha_{0},\alpha_{1},1-\alpha_{0}+\alpha_{2},3\alpha_{1}+2\alpha_{2}\},\{1-\alpha_{0},\alpha_{1},\alpha_{1}+\alpha_{2},3\alpha_{1}+2\alpha_{2}\},$ (3)
@@
This verifies the $\mathbb{F}_{q}$-stability of $\lambda$. By Theorem 3.17, fixing a nontrivial additive character $\chi:\mathbb{F}_{q}\to\mathbb{C}^{\times}$, the compactly induced representation $\pi_{\lambda}:=\mathrm{ind}_{G(k)_{x,+}}^{G(k)}\chi_{\lambda}$ is a finite direct sum of irreducible supercuspidal representations of $G(k)$.
@@
Let $n\in N(k)$ with image $w\in W(G)$ and $a,b\in G(k)_{x,0}$ such that $a^{-1}nb\in I(G,G(k)_{x,+},\chi_{\lambda})$. It suffices to show that $w(x)=x$, as then this implies $a^{-1}nb\in G(k)_{x,\lambda}$. Any lift $n_{\alpha_{2}}\in N(k)$ of the reflection in $\alpha_{2}=0$ lies in $G(k)_{x,0}=G(k)_{x}$ as it fixes $x$, where this last equality holds because $G$ is simply connected. By replacing $n$ by $nn_{\alpha_{2}}^{-1}$ and $b$ by $n_{\alpha_{2}}b$ if necessary, we may assume that $\mathrm{supp}(b\cdot\lambda)\supset\{1-\alpha_{0},\alpha_{0}\}$. Since $w(1-\alpha_{0})+w(\alpha_{0})=1$ is positive on $x$, at least one of $w(1-\alpha_{0})$ and $w(\alpha_{0})$ is in $\Psi_{x}^{+}$. By Lemma 3.18, this implies that
@@
as we have seen in (3). In each pair, a positive linear combination is $1-\alpha_{0}$ or $\alpha_{0}$, and it follows that $w$ must map at least one element $\psi$ of any pair contained in $\mathrm{supp}(b\cdot\lambda)$ to an element of $\Psi_{x}^{+}$. By Lemma 3.18, such an element must also be in $\mathrm{supp}(a\cdot\lambda)\subset\mathcal{S}(x)$. Let $\dot{w}$ denote the image of $w$ under the projection $W(G)\to W(\Phi)$. In particular, we must have that
$\dot{w}(\dot{\psi})\in\{\alpha_{1},\alpha_{1}+\alpha_{2},3\alpha_{1}+\alpha_{2},3\alpha_{1}+\alpha_{2},-\alpha_{0},-\alpha_{0}+\alpha_{2}\}.$ (4)
Suppose that $w(\alpha_{0})=1-\alpha_{0}$. Then $\dot{w}$ is either $w_{\alpha_{0}}$ or the long element of the Weyl group, $w_{0}=-1$. In the case $\dot{w}=w_{\alpha_{0}}$, in order for (4) to hold, and for $w(\psi)\in\mathcal{S}(x)$ for such $\psi$, we must have that $\psi=\alpha_{1}$ and $w(\alpha_{1})=\alpha_{1}$. Together with the equations $w(\alpha_{0})=1-\alpha_{0}$ and $\alpha_{0}=3\alpha_{1}+2\alpha_{2}$, this implies that
@@
which is impossible. In the case $\dot{w}=-1$, in order for (4) to hold, and for $w(\psi)\in\mathcal{S}(x)$ for such $\psi$, we must have that $w(1-\alpha_{0}+\alpha_{2})=3\alpha_{1}+\alpha_{2}=\alpha_{0}-\alpha_{2}$. Together with the equations $w(\alpha_{0})=1-\alpha_{0}$ and $\alpha_{0}=3\alpha_{1}+2\alpha_{2}$, this implies that $w(\alpha_{2})=-\alpha_{2}$ and so
@@
Suppose that $w(\alpha_{0})=\alpha_{0}$. Then $\dot{w}$ is either $w_{\alpha_{1}}$ or the identity element of the Weyl group. In the case $\dot{w}=w_{\alpha_{1}}$, (4) cannot occur. If $\dot{w}=1$, then by Lemma 3.18, we find that $w$ must fix an element of one of the three pairs of affine roots, so that $w$ is trivial and indeed $w(x)=x$. ∎
@@
where $\mathrm{std}$ denotes the standard representation of $\mathrm{GL}_{2}$. Suppose $g\in\mathrm{GL}_{2}(\mathbb{F}_{q})$ stabilises $\lambda$. We previously set $\lambda_{\alpha_{0}-\alpha_{2}}=0$ and $\lambda_{-\alpha_{1}}=0$, so $g$ must stabilise $\begin{pmatrix}1\\ 0\end{pmatrix}\in\mathrm{std}\otimes\det$ and $\begin{pmatrix}0\\ 1\end{pmatrix}\in\mathrm{std}\otimes\det^{-1}$. The only such matrix is the identity. We conclude that $G(k)_{x,\lambda}=G(k)_{x,+}$.
Corollary 4.9.
@@
By the proof of Proposition 3.10, if $\pi_{\lambda}$ has depth $r$, then there exists $y\in\mathcal{A}$ and $g\in\mathbb{G}_{x}(\mathbb{F}_{q})$ such that $g\cdot\lambda$ is trivial on $G(k)_{x,+}\cap G(k)_{y,r+}$. For any $g$, by (3), the support of $g\cdot\lambda$ contains $\{\psi,1-\psi\}$ for one of $\psi=\alpha_{0},\alpha_{0}-\alpha_{2}$. But for $g\cdot\lambda$ to be trivial on $G(k)_{x,+}\cap G(k)_{y,r+}$, we must then have that $\psi(y)\leq r$ and $(1-\psi)(y)\leq r$ so that $r\geq 1/2$.
On the other hand, setting $z$ to be the point with Kac coordinates $30\Rrightarrow 1$ (so $z=\frac{1}{6}\omega_{1}^{\vee}$), we have $\alpha_{0}(z)=1/2$, $(3\alpha_{1}+\alpha_{2})(z)=1/2$ and $(\alpha_{1}+\alpha_{2})(z)=1/6\leq 1/2$. Thus $\lambda$ is trivial on $G(k)_{z,\frac{1}{2}+}$ and so $\pi_{\lambda}$ has a nonzero fixed vector under $G(k)_{z,\frac{1}{2}+}$. Therefore, $\pi_{\lambda}$ is epipelagic of depth $1/2$. ∎
Proposition 4.11.

### v2

@@ 4.1 $G=G_{2}$, $q=2$
As $G(k)_{x,+}=G(k)_{x,0}$, it suffices to check that the only element $w$ of the affine Weyl group that intertwines $\chi_{\lambda_{i}}$ is trivial. By Lemma 3.3.8,
@@
and so $w$ is either trivial or the simple reflection $w_{\alpha_{1}}$ in $\alpha_{1}=0$. The latter sends $\alpha_{2}$ to $3\alpha_{1}+\alpha_{2}\in\Psi_{x}^{+}$, violating the condition from Lemma 3.3.8. Therefore $w$ must be trivial. Lemma 3.1.2 implies that the $\pi_{i}$ are irreducible and supercuspidal for $i=1,2$. ∎
@@ 4.2 $G=B_{n}$, $q=2$
These inequalities imply $r_{i}=0$ for all $i$, as required. Then, by Theorem 3.3.7, letting $\chi$ be the unique nontrivial additive character $\chi:\mathbb{F}_{2}\to\mathbb{C}^{\times}$, the compactly induced representation $\pi_{\lambda}:=\mathrm{ind}_{G(k)_{x,+}}^{G(k)}\chi_{\lambda}$ is a finite direct sum of irreducible supercuspidal representations of $G(k)$.
Remark 4.2.1.
@@
By the proof of Proposition 3.2.6, if $\pi_{\lambda}$ has a nontrivial fixed vector under $G(k)_{y^{\prime},r+}$ for some $y^{\prime}\in\mathcal{B}$, then there exists $y\in\mathcal{A}$ and $g\in\mathbb{G}_{x}(\mathbb{F}_{q})$ such that $g\cdot\lambda$ is trivial on $G(k)_{x,+}\cap G(k)_{y,r+}$. In our case $\mathbb{G}_{x}(\mathbb{F}_{2})$ is trivial. Let $y=\sum_{i}r_{i}\omega_{i}^{\vee}$ for $r_{i}\in\mathbb{R}$. Then $\lambda$ is trivial on $G(k)_{x,+}\cap G(k)_{y,r+}$ if and only if the following inequalities hold:
@@
Combining [RY14, Theorem 4.1] and [RLYG12, Section 8.2.2], the points $x\in\mathcal{A}$ at which $V_{x}^{\vee}$ admits an $\mathbb{F}_{q}$-rational stable functional satisfy $\delta(x)=m/2n$, where $m\mid n$. By Proposition 3.2.6, the corresponding supercuspidal representations constructed by [RY14, Proposition 2.4] have depth $m/2n$. When $n\geq 3$ this does not take the value $\frac{1}{2(n-1)}$. ∎
@@
As we are working in the setting $q=2$, we have that $G(k)_{x,+}=G(k)_{x,0}$. Note that in the simply-connected case, the $G(k)_{x,+}$-double cosets of $1$ and $\sigma$ agree, and also of $w_{\alpha_{n}}$ and $\sigma w_{\alpha_{n}}$ agree. By the affine Bruhat decomposition 3.1.4, it suffices to show that the elements of $W(G)$ that intertwine $\chi_{\lambda}$ are $1,w_{\alpha_{n}},\sigma$ and $\sigma w_{\alpha_{n}}$. These certainly intertwine as their action on $\Psi$ takes $\mathrm{supp}(\lambda)$ to itself so that $\lambda=\lambda^{w}$ on $G(k)_{x,+}\cap G(k)_{x,+}^{w}$ for any of $w=1,w_{\alpha_{n}},\sigma$ and $\sigma w_{\alpha_{n}}$.
On the other hand, by Lemma 3.3.8, if $w\in W(G)$ intertwines $\chi_{\lambda}$, then
@@
Moreover, by Lemma 3.3.9, the set $w(\mathcal{S}(x))\cap\mathcal{S}(x)$ has no zeros in $\overline{\mathcal{C}}$. This set must therefore contain $1-\alpha_{0},\alpha_{1},\dots,\alpha_{n-2}$ and also $\{\alpha_{n-1}+\alpha_{n}\}$, $\{\alpha_{n-1},\alpha_{n}\}$, $\{\alpha_{n-1},\alpha_{n-1}+2\alpha_{n}\}$ or $\{\alpha_{n},\alpha_{n-1}+2\alpha_{n}\}$. The only two elements of $\mathcal{S}(x)$ whose gradients pair negatively with the gradients of at least three other elements of $\mathcal{S}(x)$ are $\alpha_{2}$ and $\alpha_{n-2}$. For $\alpha_{2}$, these three elements are $1-\alpha_{0},\alpha_{1}$ and $\alpha_{3}$. As these are all in $w(\mathcal{S}(x))\cap\mathcal{S}(x)$ (as $n\geq 5$), it follows that $w(\alpha_{2})=\alpha_{2}$ or $w(\alpha_{n-2})=\alpha_{2}$. Here we use the fact that the Weyl group acts by isometries on the roots.
@@
Given $w(\alpha_{2})=\alpha_{2}$, again using combinatorics of the affine Dynkin diagram, $w$ must fix $\alpha_{3},\dots,\alpha_{n-2}$ and either fix or swap $1-\alpha_{0},\alpha_{1}$. Composing with $\sigma$, which swaps $1-\alpha_{0},\alpha_{1}$, we may assume that $w$ fixes them. We must show that $w=1$ or $w=w_{\alpha_{n}}$. Out of the remaining elements $\alpha_{n-1}$, $\alpha_{n}$, $\alpha_{n-1}+\alpha_{n}$ and $\alpha_{n-1}=2\alpha_{n}$, the only one to pair with $\alpha_{n-2}$ to 0 is $\alpha_{n}$. So if $\alpha_{n}\in w(\mathcal{S}(x))\cap\mathcal{S}(x)$ then $w(\alpha_{n})=\alpha_{n}$ and the only possibility is $w=1$. Similarly, $\alpha_{n-1}+\alpha_{n}$ is the only short root of the four that pairs negatively with $\alpha_{n-2}$. If $\alpha_{n-1}+\alpha_{n}\in w(\mathcal{S}(x))\cap\mathcal{S}(x)$ then $w(\alpha_{n-1}+\alpha_{n})=\alpha_{n-1}+\alpha_{n}$. One of $\alpha_{n-1}$ or $\alpha_{n-1}+2\alpha_{n}$ must then map under $w$ to a positive affine root at $x$, necessarily in $\mathcal{S}(x)$ by Lemma 3.3.8. The only possibility is that
@@
Assume now that $G=\mathrm{Spin}_{2n+1}$ with $n\geq 5$. We will explicitly describe the two irreducible supercuspidal representations above as compactly induced representations. As in Proposition 4.2.2, take $y\in\overline{\mathcal{C}}$ to be the point with $\alpha_{1}(y)=\dots=\alpha_{n-1}(y)=\frac{1}{2(n-1)}$ and $\alpha_{n}(y)=0$. Then
$\Delta(y)=\{1-\alpha_{0},\alpha_{1},\dots,\alpha_{n-1},\alpha_{n-1}+\alpha_{n},\alpha_{n-1}+2\alpha_{n}\}.$
Over $\mathbb{F}_{2}$, noting that $\mathbb{G}_{m}(\mathbb{F}_{2})=1$, the reductive quotient at $y$ has $\mathbb{F}_{2}$-points $\mathbb{G}_{y}(\mathbb{F}_{2})=\mathrm{SL}_{2}(\mathbb{F}_{2})$. Let $\lambda^{\prime}$ denote the $\mathbb{F}_{2}$-linear functional on $G(k)_{y,+:++}$ with support $\Delta(y)$. This is not stable for the action of $\mathbb{G}_{y}$, as we know from the classification of stable points in [RLYG12] (see Corollary 4.2.3). However, it is $\mathbb{F}_{2}$-stable as $\mathbb{G}_{y}(\mathbb{F}_{2})=\mathrm{SL}_{2}(\mathbb{F}_{2})$ acts trivially on $\lambda^{\prime}$. The argument of [RY14] continues to imply that the intertwining set satisfies $I(G,G(k)_{y,+},\chi_{\lambda^{\prime}})\subset G(k)_{y}=G(k)_{y,0}$. This could otherwise be seen using the same argument as Proposition 4.2.4, where this time any lift to $N(k)$ of $w_{\alpha_{n}}$ is in $G(k)_{y}$. In fact we have
$I(G,G(k)_{y,+},\chi_{\lambda^{\prime}})=G(k)_{y}$
because $\mathbb{G}_{y}(\mathbb{F}_{2})$ acts trivially on $\lambda^{\prime}$. By [RY14, Lemma 2.2, Remark 1], the representation $\mathrm{ind}_{G(k)_{y,+}}^{G(k)}\chi_{\lambda^{\prime}}$ decomposes as a direct sum of three non-isomorphic supercuspidals, indexed by the three irreducible representations of $\mathrm{SL}_{2}(\mathbb{F}_{2})\cong S_{3}$. The image of the Iwahori subgroup $G(k)_{x}=G(k)_{x,+}\subset G(k)_{y}$ in $\mathbb{G}_{y}(\mathbb{F}_{2})$ is a subgroup of order 2, given by $U_{\alpha_{n}}/U_{\alpha_{n}+1}$. The two nontrivial irreducible representations $\rho_{1},\rho_{2}$ of $\mathrm{SL}_{2}(\mathbb{F}_{2})$ are nontrivial on any such order 2 subgroup. In the notation of Section 3.1, it follows that
$\mathrm{Hom}_{G(k)}\left(\mathrm{ind}_{G(k)_{x}}^{G(k)}\chi_{\lambda},\mathrm{ind}_{G(k)_{y}}^{G(k)}(\chi_{\lambda^{\prime}})_{\rho_{i}}\right)\neq 0$
for $i=1,2$, because $\lambda$ is nontrivial on $U_{\alpha_{n}}/U_{\alpha_{n}+1}$. Therefore
$\pi_{\lambda}\cong\mathrm{ind}_{G(k)_{y}}^{G(k)}(\chi_{\lambda^{\prime}})_{\rho_{1}}\oplus\mathrm{ind}_{G(k)_{y}}^{G(k)}(\chi_{\lambda^{\prime}})_{\rho_{2}}.$
The proof of Proposition 3.2.6 applies here to show that both of these irreducible supercuspidals are epipelagic of depth $\frac{1}{2(n-1)}$.
@@ 4.3 $G=G_{2}$, $p=3$
with $\lambda_{-\dot{\psi}}\in V_{\psi}^{\vee}$. Set $\lambda_{-(\alpha_{1}+\alpha_{2})}=0,\lambda_{-(3\alpha_{1}+\alpha_{2})}=0$. Take $\lambda_{\alpha_{0}}$, $\lambda_{\alpha_{0}-\alpha_{2}}$, $\lambda_{-\alpha_{1}}$ and $\lambda_{-(3\alpha_{1}+2\alpha_{2})}$ to be any nonzero $\mathbb{F}_{q}$-linear functionals in the respective $V_{\psi}^{\vee}$. We claim that this gives an $\mathbb{F}_{q}$-stable functional.
Let $g\in\mathbb{G}_{x}(\mathbb{F}_{q})$. The factor $\mathbb{G}_{m}$ of $\mathrm{SL}_{2}\times\mathbb{G}_{m}$ acts by nonzero scaling on each summand of $V(x)^{\vee}$. Identify $V_{1-\alpha_{0}}^{\vee}\oplus V_{1-\alpha_{0}+\alpha_{2}}^{\vee}$, $V_{\alpha_{1}}^{\vee}\oplus V_{\alpha_{1}+\alpha_{2}}^{\vee}$ and $V_{3\alpha_{1}+\alpha_{2}}^{\vee}\oplus V_{3\alpha_{1}+2\alpha_{2}}^{\vee}$ as the standard representation of $\mathrm{SL}_{2}$. As $(\lambda_{\alpha_{0}},\lambda_{\alpha_{0}-\alpha_{2}})$, $(\lambda_{-\alpha_{1}},0)$ and $(0,\lambda_{-(3\alpha_{1}+2\alpha_{2})})$ are pairwise linearly independent under this identification, so are their images under the action of $g$. Therefore, $\mathrm{supp}(g\cdot\lambda)$ contains at least two elements of each of $\{1-\alpha_{0},\alpha_{1},3\alpha_{1}+\alpha_{2}\}$ and $\{1-\alpha_{0}+\alpha_{2},\alpha_{1}+\alpha_{2},3\alpha_{1}+2\alpha_{2}\}$. It also contains at least one element of each of $\{1-\alpha_{0},1-\alpha_{0}+\alpha_{2}\}$, $\{\alpha_{1},\alpha_{1}+\alpha_{2}\}$ and $\{3\alpha_{1}+\alpha_{2},3\alpha_{1}+2\alpha_{2}\}$.
@@
$\{1-\alpha_{0},\alpha_{1},1-\alpha_{0}+\alpha_{2},3\alpha_{1}+2\alpha_{2}\},\{1-\alpha_{0},\alpha_{1},\alpha_{1}+\alpha_{2},3\alpha_{1}+2\alpha_{2}\},$ (4.3.1)
@@
This verifies the $\mathbb{F}_{q}$-stability of $\lambda$. By Theorem 3.3.7, fixing a nontrivial additive character $\chi:\mathbb{F}_{q}\to\mathbb{C}^{\times}$, the compactly induced representation $\pi_{\lambda}:=\mathrm{ind}_{G(k)_{x,+}}^{G(k)}\chi_{\lambda}$ is a finite direct sum of irreducible supercuspidal representations of $G(k)$.
@@
Let $n\in N(k)$ with image $w\in W(G)$ and $a,b\in G(k)_{x,0}$ such that $a^{-1}nb\in I(G,G(k)_{x,+},\chi_{\lambda})$. It suffices to show that $w(x)=x$, as then this implies $a^{-1}nb\in G(k)_{x,\lambda}$. Any lift $n_{\alpha_{2}}\in N(k)$ of the reflection in $\alpha_{2}=0$ lies in $G(k)_{x,0}=G(k)_{x}$ as it fixes $x$, where this last equality holds because $G$ is simply connected. By replacing $n$ by $nn_{\alpha_{2}}^{-1}$ and $b$ by $n_{\alpha_{2}}b$ if necessary, we may assume that $\mathrm{supp}(b\cdot\lambda)\supset\{1-\alpha_{0},\alpha_{0}\}$. Since $w(1-\alpha_{0})+w(\alpha_{0})=1$ is positive on $x$, at least one of $w(1-\alpha_{0})$ and $w(\alpha_{0})$ is in $\Psi_{x}^{+}$. By Lemma 3.3.8, this implies that
@@
as we have seen in (4.3.1). In each pair, a positive linear combination is $1-\alpha_{0}$ or $\alpha_{0}$, and it follows that $w$ must map at least one element $\psi$ of any pair contained in $\mathrm{supp}(b\cdot\lambda)$ to an element of $\Psi_{x}^{+}$. By Lemma 3.3.8, such an element must also be in $\mathrm{supp}(a\cdot\lambda)\subset\mathcal{S}(x)$. Let $\dot{w}$ denote the image of $w$ under the projection $W(G)\to W(\Phi)$. In particular, we must have that
$\dot{w}(\dot{\psi})\in\{\alpha_{1},\alpha_{1}+\alpha_{2},3\alpha_{1}+\alpha_{2},3\alpha_{1}+\alpha_{2},-\alpha_{0},-\alpha_{0}+\alpha_{2}\}.$ (4.3.3)
Suppose that $w(\alpha_{0})=1-\alpha_{0}$. Then $\dot{w}$ is either $w_{\alpha_{0}}$ or the long element of the Weyl group, $w_{0}=-1$. In the case $\dot{w}=w_{\alpha_{0}}$, in order for (4.3.3) to hold, and for $w(\psi)\in\mathcal{S}(x)$ for such $\psi$, we must have that $\psi=\alpha_{1}$ and $w(\alpha_{1})=\alpha_{1}$. Together with the equations $w(\alpha_{0})=1-\alpha_{0}$ and $\alpha_{0}=3\alpha_{1}+2\alpha_{2}$, this implies that
@@
which is impossible. In the case $\dot{w}=-1$, in order for (4.3.3) to hold, and for $w(\psi)\in\mathcal{S}(x)$ for such $\psi$, we must have that $w(1-\alpha_{0}+\alpha_{2})=3\alpha_{1}+\alpha_{2}=\alpha_{0}-\alpha_{2}$. Together with the equations $w(\alpha_{0})=1-\alpha_{0}$ and $\alpha_{0}=3\alpha_{1}+2\alpha_{2}$, this implies that $w(\alpha_{2})=-\alpha_{2}$ and so
@@
Suppose that $w(\alpha_{0})=\alpha_{0}$. Then $\dot{w}$ is either $w_{\alpha_{1}}$ or the identity element of the Weyl group. In the case $\dot{w}=w_{\alpha_{1}}$, (4.3.3) cannot occur. If $\dot{w}=1$, then by Lemma 3.3.8, we find that $w$ must fix an element of one of the three pairs of affine roots, so that $w$ is trivial and indeed $w(x)=x$. ∎
@@
where $\mathrm{std}$ denotes the standard representation of $\mathrm{GL}_{2}$. Suppose $g\in\mathrm{GL}_{2}(\mathbb{F}_{q})$ stabilises $\lambda$. We previously set $\lambda_{-(\alpha_{1}+\alpha_{2})}=0$ and $\lambda_{-(3\alpha_{1}+\alpha_{2})}=0$, so $g$ must stabilise $\begin{pmatrix}1\\ 0\end{pmatrix}\in\mathrm{std}\otimes\det^{-1}$ and $\begin{pmatrix}0\\ 1\end{pmatrix}\in\mathrm{std}\otimes\det^{-2}$. The only such matrix is the identity. We conclude that $G(k)_{x,\lambda}=G(k)_{x,+}$.
Corollary 4.3.4.
@@
By the proof of Proposition 3.2.6, if $\pi_{\lambda}$ has depth $r$, then there exists $y\in\mathcal{A}$ and $g\in\mathbb{G}_{x}(\mathbb{F}_{q})$ such that $g\cdot\lambda$ is trivial on $G(k)_{x,+}\cap G(k)_{y,r+}$. For any $g$, by (4.3.1), the support of $g\cdot\lambda$ contains $\{\psi,1-\psi\}$ for one of $\psi=\alpha_{0},\alpha_{0}-\alpha_{2}$. But for $g\cdot\lambda$ to be trivial on $G(k)_{x,+}\cap G(k)_{y,r+}$, we must then have that $\psi(y)\leq r$ and $(1-\psi)(y)\leq r$ so that $r\geq 1/2$.
On the other hand, setting $z=\frac{1}{2}(\omega_{1}^{\vee}-\omega_{2}^{\vee})$, we have $1-\alpha_{0}(z)=\alpha_{0}(z)=1/2$, $\alpha_{1}(z)=1/2$ and $1-(3\alpha_{1}+\alpha_{2})(z)=0\leq 1/2$. Thus $\lambda$ is trivial on $G(k)_{z,\frac{1}{2}+}$ and so $\pi_{\lambda}$ has a nonzero fixed vector under $G(k)_{z,\frac{1}{2}+}$. Therefore, $\pi_{\lambda}$ is epipelagic of depth $1/2$. ∎
Proposition 4.3.6.